If the pressure amplitude of a sound wave is doubled, then intensity of sound will increase to a factor of

Grade:11

1 Answers

Manasa S
6 years ago
I=2π²vf²A²(rho)I directly proportional A^2I/I`=(A/A`)²A`=2AI:I`=1:4I`=4IIntensity increases to the factor 4